4 Use the following data in the table to answer the Amount of Buils Bridges Bricks questions a. Independent variable: b. Dependent variable: c. Domains minimum to d· The rate at which the amount of bricks 10 20 30 40 50 60 70 12 15 18 21 makes the number of bridges e. Graph the data. Label the graph. 5. You are buying a new Honda CRZ Hybrid for $28,650. You make a down payment of $2750 a. How much do you need to borrow?ーーーーーーーーーーーー b. you take out a loan for 6 years at 5.5%, what is your monthly payment? APR PMT = APR-nt
Answer #1

4 a )  

independent variable = amount of bricks

b) dependent variable = number of bridges

c) domain minimum = 10 to maximum = 70

d) the rate at which amount of bricks maks the number of bridges = (6-3) / ( 20 - 10 ) = 3/10

e) graph shown below

30 20 10 30 40 50 60 70 10 20